# Strange metal and superconductor in the two-dimensional Yukawa-Sachdev-Ye-Kitaev model

### v2

@@ Abstract
Our paper presents a self-consistent, disorder-averaged analysis of a two-dimensional Yukawa-Sachdev-Ye-Kitaev (2d-YSYK) model, which has a spatially random Yukawa coupling between fermions, $\psi$, with a Fermi surface and a nearly-critical scalar field, $\phi$. We use methods similar to those which yield the exact solution of the zero-dimensional Sachdev-Ye-Kitaev model. Such a 2d-YSYK model has been argued [7, 8, 9] to provide a universal description of quantum phase transitions in metals, associated with the condensation of $\phi$, in the presence of impurity-induced ‘Harris’ disorder [10, 11, 12] with spatial fluctuations in the local position of the quantum critical point. We find results that display all the key characteristics of the optical conductivity and d.c. resistivity described by Michon et al., as shown in Fig. 3.
@@
$\mathcal{L}_{\rm 2d-YSYK}=\psi^{\dagger}_{\sigma}\left[\partial_{\tau}+\varepsilon(i\nabla)-\mu\right]\psi_{\sigma}+v({\bm{r}})\psi^{\dagger}_{\sigma}\psi_{\sigma}\,$
$~{}~{}+\frac{1}{2}\left[(\partial_{\tau}\phi)^{2}+c^{2}(\nabla\phi)^{2}+s\phi^{2}\right]+u\phi^{4}$
$~{}~{}+[g+g^{\prime}({\bm{r}})]\,\phi_{\sigma}\psi_{\sigma}^{\dagger}\mathcal{D}_{\bm{r}}\psi_{\sigma}.$ (1)
Here $\tau$ is imaginary time, we set $\hbar=1$, $s$ is the tuning parameter across the transition, $u$ is a scalar self-interaction. The operator ${\mathcal{D}}_{\bm{r}}=\partial_{x}^{2}-\partial_{y}^{2}$ is special to the Ising-nematic case, and will be set to unity for simplicity in our computations as it unimportant apart from ‘cold spots’ on the Fermi surface. In order to obtain a spin-singlet superconductor we have spinful fermions, with $\sigma$ the spin index.
$\mathcal{L}_{\rm 2d-YSYK}$ contains the two sources of disorder. One is the potential $v({\bm{r}})$ acting on the fermions:
  • Spatially random potential $v({\bm{r}})$ with $\overline{v({\bm{r}})}=0$, $\overline{v({\bm{r}})v({\bm{r}^{\prime}})}=v^{2}\delta({\bm{r}}-{\bm{r}^{\prime}})$.
Its influence is familiar from the theory of weakly disordered metals [35], leading to marginally relevant localization effects on the fermions [35]. Much more relevant is the Harris disorder, which we have taken in the form of a spatially random Yukawa coupling $g^{\prime}({\bm{r}})$ adding to the spatially uniform Yukawa coupling $g$, and obeying
@@
By rescaling $\phi$, it is possible to transform $g^{\prime}({\bm{r}})$ into a more familiar ‘random mass’ form of the Harris disorder, $s\rightarrow s+\delta s({\bm{r}})$ [7]. Both forms of the disorder have been examined in earlier work, random $g^{\prime}({\bm{r}})$ [7, 9] and random $\delta s({\bm{r}})$ [8], and similar results were obtained. We choose to work in the random $g^{\prime}({\bm{r}})$ formulation because it enables direct extension of methods employed for the exact solution of the zero-dimensional YSYK model [36, 37, 38, 39, 13, 14, 40, 15, 41]. We restrict our analysis here to the simpler case with $g=0$ for then the self-energies are functions of frequency alone; earlier work [7] has shown that perturbative corrections in $g$ cancel in the transport response, partially justifying the $g=0$ choice.
Two distinct regimes of behavior of $\mathcal{L}_{\rm 2d-YSYK}$ have been identified [8, 9]:
(i) There is a significant intermediate energy regime where the bosonic and fermionic eigenmodes are spatially extended. In this regime, the physics is self-averaging, and numerical results are consistent with the averaged Green’s function methods which yield the exact solution of the zero-dimensional YSYK model. This is the regime we address in the present paper by standard SYK methods.
(ii) At low $T$ there is a crossover to a regime where bosonic eigenmodes are localized, while the fermionic eigenmodes remain extended [42]. Here, we must treat the disorder in the bosonic sector carefully: by a strong disorder renormalization [12], numerically exactly [8, 9], or map to ‘two-level system’ models [43, 44]. Recent analyses [45, 46] of $v({\bm{r}})$ disorder effects along the lines of Ref. 35 near quantum criticality found singular corrections to the boson propagator, which we view as a precursor to boson localization. This boson localization regime is not described in the present paper.
@@
where the density of states $\rho_{{\rm tr}}^{\alpha\beta}\left(\epsilon\right)\equiv\int_{\bm{k}}\frac{\partial\varepsilon_{\bm{k}}}{\partial k_{\alpha}}\frac{\partial\varepsilon_{\bm{k}}}{\partial k_{\beta}}\delta\left(\epsilon-\varepsilon_{\bm{k}}\right)$. In Fig. 5, we show the temperature dependence of $\rho_{s}$ for $g^{\prime}=5t^{3/2}$ as the system is tuned away from the QCP by varying the boson mass $M$. The stiffness reaches a finite constant value, $\rho_{s0}$, for $T\rightarrow 0$. As the system is tuned away from the QCP, $T_{c}$ decreases from its maximal value while $\rho_{s0}$ increases, as seen in Fig. 5. This result at strong coupling aligns with the superfluid density in bulk samples of the overdoped cuprate superconductors [50, 19]. When $T_{c}$ is largest at the QCP, the opening of a $\psi$ gap at $T=0$ weakens the pairing interaction mediated by $\phi$ (because the $\phi$ self energy $\Pi$ is self-consistently determined by the $\psi$ polarization, see Fig. 1), leading to a smaller superfluid stiffness at $T=0$ [16, 17]. On the other hand, in the presence of strong potential scattering $v$ this effect is weaker, and a decreasing $T_{c}$ is eventually accompanied by a decreasing $\rho_{s0}$, as shown in Fig. LABEL:fig:stiffness_s [21].

### v3

@@ Abstract
Our paper presents a self-consistent, disorder-averaged analysis of a two-dimensional Yukawa-Sachdev-Ye-Kitaev (2d-YSYK) model, which has a spatially random Yukawa coupling between fermions, $\psi$, with a Fermi surface and a nearly-critical scalar field, $\phi$. We use methods similar to those which yield the exact solution of the zero-dimensional Sachdev-Ye-Kitaev model. Such a 2d-YSYK model has been argued [7, 8, 9, 10] to provide a universal description of quantum phase transitions in metals, associated with the condensation of $\phi$, in the presence of impurity-induced ‘Harris’ disorder [11, 12, 13] with spatial fluctuations in the local position of the quantum critical point. We find results that display all the key characteristics of the optical conductivity and d.c. resistivity described by Michon et al., as shown in Fig. 3.
@@
$\mathcal{L}_{\rm 2d-YSYK}=\psi^{\dagger}_{i\sigma}\left[\partial_{\tau}+\varepsilon(i\nabla)-\mu\right]\psi_{i\sigma}+\frac{v_{ij}}{\sqrt{N}}({\bm{r}})\psi^{\dagger}_{i\sigma}\psi_{i\sigma}\,$
$~{}~{}+\frac{1}{2}\left[(\partial_{\tau}\phi_{i})^{2}+c^{2}(\nabla\phi_{i})^{2}+s\phi_{i}^{2}\right]+u\phi_{i}^{4}$
$~{}~{}+\frac{1}{N}[g_{ij\ell}+g^{\prime}_{ij\ell}({\bm{r}})]\,\phi_{i\ell}\psi_{i\sigma}^{\dagger}\mathcal{D}_{\bm{r}}\psi_{j\sigma}\,.$ (1)
The flavor indices $i,j,\ell$ are summed over $N$ values—the equations in Fig. 1 are exact at all $T$ in the large $N$ limit when the couplings are random in flavor space. Our discussion is for the physical case $N=1$, for which Fig. 1 is applied over an intermediate $T$ regime. Imaginary time is $\tau$, we set $\hbar=1$, $s$ is the tuning parameter across the transition, $u$ is a scalar self-interaction. The operator ${\mathcal{D}}_{\bm{r}}=\partial_{x}^{2}-\partial_{y}^{2}$ is special to the Ising-nematic case, and will be set to unity for simplicity in our computations as it unimportant apart from ‘cold spots’ on the Fermi surface. In order to obtain a spin-singlet superconductor we have spinful fermions, with $\sigma$ the spin index.
$\mathcal{L}_{\rm 2d-YSYK}$ contains the two sources of disorder. One is the potential $v({\bm{r}})$ acting on the fermions. It is random in both flavor and space (in (2) and (3) we omit flavor indices for clarity, see the supplement [22] for full forms):
$\overline{v({\bm{r}})}=0,\quad\overline{v({\bm{r}})v({\bm{r}^{\prime}})}=v^{2}\delta({\bm{r}}-{\bm{r}^{\prime}}).$ (2)
Its influence is familiar from the theory of weakly disordered metals, leading to marginally relevant localization effects on the fermions [35]. Much more relevant is the Harris disorder, which we have taken in the form of a spatially random Yukawa coupling $g^{\prime}({\bm{r}})$ adding to the spatially uniform Yukawa coupling $g$, and obeying
@@
By rescaling $\phi$, it is possible to transform $g^{\prime}({\bm{r}})$ into a more familiar ‘random mass’ form of the Harris disorder, $s\rightarrow s+\delta s({\bm{r}})$ [8]. Both forms of the disorder have been examined in earlier work, random $g^{\prime}({\bm{r}})$ [8, 10] and random $\delta s({\bm{r}})$ [9], and similar results were obtained. We choose to work in the random $g^{\prime}({\bm{r}})$ formulation because it enables direct extension of methods employed for the exact solution of the zero-dimensional YSYK model [36, 37, 38, 39, 14, 15, 40, 16, 41]. We restrict our analysis here to the simpler case with $g=0$ for then the self-energies are functions of frequency alone; earlier work [8] has shown that perturbative corrections in $g$ cancel in the transport response, partially justifying the $g=0$ choice. Two distinct regimes of behavior of $\mathcal{L}_{\rm 2d-YSYK}$ have been identified [9, 10]:
(i) There is a significant intermediate energy regime where the bosonic and fermionic eigenmodes are spatially extended. In this regime, the physics is self-averaging, and numerical results are consistent with the large-$N$ averaged Green’s function methods which yield the exact solution of the zero-dimensional YSYK model. This is the regime we address in the present paper by standard SYK methods. The spatial disorder is needed for the singular single-particle self-energies to feed into singular transport properties [8], which does not happen for translationally invariant models [23, 24, 25, 26, 27, 28, 29, 30].
(ii) At low $T$ there is a crossover to a regime where spatial disorder causes bosonic eigenmodes to localize [9, 10], while the fermionic eigenmodes remain extended [42]. Here, we must treat the disorder in the bosonic sector more completely: by a strong disorder renormalization [13], numerically exactly [9, 10], or map to ‘two-level system’ models [43, 44]. Recent analyses [45, 46] of $v({\bm{r}})$ disorder effects along the lines of Ref. 35 near quantum criticality found singular corrections to the boson propagator, which we view as a precursor to boson localization. This boson localization regime is not described in the present paper.
@@
where the density of states $\rho_{{\rm tr}}^{\alpha\beta}\left(\epsilon\right)\equiv\int_{\bm{k}}\frac{\partial\varepsilon_{\bm{k}}}{\partial k_{\alpha}}\frac{\partial\varepsilon_{\bm{k}}}{\partial k_{\beta}}\delta\left(\epsilon-\varepsilon_{\bm{k}}\right)$. In Fig. 5, we show the temperature dependence of $\rho_{s}$ for $g^{\prime}=5t^{3/2}$ as the system is tuned away from the QCP by varying the boson mass $M$. The stiffness reaches a finite constant value, $\rho_{s0}$, for $T\rightarrow 0$. As the system is tuned away from the QCP, $T_{c}$ decreases from its maximal value while $\rho_{s0}$ increases, as seen in Fig. 5. This result at strong coupling aligns with the superfluid density in bulk samples of the overdoped cuprate superconductors [50, 20]. When $T_{c}$ is largest at the QCP, the opening of a $\psi$ gap at $T=0$ weakens the pairing interaction mediated by $\phi$ (because the $\phi$ self energy $\Pi$ is self-consistently determined by the $\psi$ polarization, see Fig. 1), leading to a smaller superfluid stiffness at $T=0$ [17, 18]. On the other hand, in the presence of strong potential scattering $v$ this effect is weaker, and a decreasing $T_{c}$ is eventually accompanied by a decreasing $\rho_{s0}$, as shown in Fig. LABEL:fig:stiffness_s [22]. The low $T$ boson localization [9] could have a significant influence on the spatial inhomogeneity of the superfluid density, and this remains to be studied.